# An Autonomous Vision-Based Algorithm for Interplanetary Navigation

### v2

@@ 3.2 Light-Aberration Correction
After the first attitude determination, the centroids of the stars are corrected for the light-aberration effect, and the probe attitude is recomputed by taking into account the corrected stars LoS directions. The procedure adopted is described in [19] and summarized in this section. At first, the observed stars LoS directions as seen by the spacecraft in the inertial reference frame $\mathcal{I}$ are found:
${}^{\mathcal{I}}\bm{\rho}_{s_{\rm{obs}}}=(\bm{K}_{\rm{cam}}\bm{A})^{-1}{}_{h}^{\mathbb{C}}\bm{r}_{s_{\rm{obs}}}$ (4)
where ${}_{h}^{\mathbb{C}}\bm{r}_{s_{\rm{obs}}}=[^{\mathbb{C}}\bm{r}_{s_{\rm{obs}}};1]^{\top}$ are the observed stars projection in $\mathbb{C}$ in homogeneous coordinates (see Hartley and Zisserman [39] for homogeneous coordinates definition). Conversely, $\bm{A}$ is the attitude matrix, and $\bm{K}_{\rm{cam}}$ is the camera calibration matrix defined as:
$\bm{K}_{\rm{cam}}=\begin{bmatrix}f_{1}&&0&&c_{1}\\ 0&&f_{2}&&c_{2}\\ 0&&0&&1\\ \end{bmatrix}$ (5)
where $f_{1}$ and $f_{2}$ represent the camera focal length along $\bm{C}_{1}$ and $\bm{C}_{2}$ directions in pixel. Whereas, $c_{1}$ and $c_{2}$ describe the position of the principal point in pixels in $\mathbb{C}$. Then, the angle $\theta_{\textrm{obs}}$ between ${}^{\mathcal{I}}\hat{\bm{\rho}}_{s_{\rm{obs}}}$ (i.e., the unit vector of ${}^{\mathcal{I}}\bm{\rho}_{s_{\rm{obs}}}$) and the estimated unitary velocity vector of the probe $\hat{\bm{v}}$ is defined as
$\tan{\theta_{\textrm{obs}}}=\frac{||{}^{\mathcal{I}}\hat{\bm{\rho}}_{s_{\rm{obs}}}\times\hat{\bm{v}}||}{{}^{\mathcal{I}}\hat{\bm{\rho}}_{s_{\rm{obs}}}^{\top}\hat{\bm{v}}}$ (6)
Then the aberration angle $\varepsilon$ is evaluated:
$\tan{\varepsilon}=\dfrac{(v/c)\sin\theta_{\textrm{obs}}}{1-(v/c)\cos\theta_{\textrm{obs}}}$ (7)
where $c$ is the speed of light. Thus, the corrected unitary stars LoS directions ${}^{\mathcal{I}}\hat{\bm{\rho}}_{s_{\rm{corr}}}$ can be retrieved as
${}^{\mathcal{I}}\hat{\bm{\rho}}_{s_{\rm{corr}}}=\dfrac{{}^{\mathcal{I}}\hat{\bm{\rho}}_{s_{\textrm{obs}}}\sin\theta_{\textrm{corr}}-\hat{\bm{v}}\sin\varepsilon}{\sin{\theta_{\textrm{obs}}}}$ (8)
with $\theta_{\textrm{corr}}=\theta_{\textrm{obs}}+\varepsilon$. At this point, the attitude matrix of the probe is redetermined by solving Wahba’s problem [37] in which ${}^{\mathcal{I}}\hat{\bm{\rho}}_{s_{\rm{corr}}}$ are considered. This corrected attitude matrix is labeled $\bm{A}_{\rm{corr}}$.
@@ 3.3 Beacon Identification
At this step, the planet must be identified in the image and its projection ${}^{\mathbb{C}}\bm{r}_{\rm{pl}}$ must be extracted. The identification is performed through the evaluation of the mean and covariance matrix associated with the planet location. This latter defines the Gaussian probability of finding the planet in that portion of the image. At first, the expected location of the observed planet in the image plane is evaluated as:
${}_{h}^{\mathbb{C}}\bm{r}_{\rm{pl}_{0}}=\bm{K}_{\rm{cam}}\bm{A}_{\rm{corr}}(^{\mathcal{I}}\bm{r}_{\rm{pl}}-^{\mathcal{I}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{r}})$ (9)
@@
$\bm{G}$ is the matrix describing the mapping between ${}^{\mathbb{C}}\bm{r}_{\rm{pl}}$ and the spacecraft pose and beacon position. $\bm{S}$ is the covariance matrix of the probe pose and beacon position. To evaluate $\bm{G}$, the variation of $\ {}^{\mathbb{C}}\bm{r}_{\rm{pl}}$ with respect to the variation of the spacecraft pose and the beacon position is computed. In particular, the quaternions $\bm{q}_{\mathcal{C}/\mathcal{N}}=(q_{0},\bm{q}_{v})^{\top}$ are chosen to represent the probe attitude matrix. Specifically, $q_{0}$ and $\bm{q}_{v}$ represent the quaternion scalar and vectorial part, respectively. Eq. (11) gives the quaternion representation of the attitude matrix $\bm{A}_{\textrm{corr}}$ [37]
@@
where $\left[{(\cdot)}\right]^{\wedge}$ is the skew-symmetric matrix associated with the cross-product operation. Thus, the variation of $\ {}^{\mathbb{C}}\bm{r}_{\rm{pl}}$ with respect to the variation of the spacecraft pose (i.e., $\bm{A}_{\textrm{corr}}$ and ${}^{\mathcal{I}}\bm{r}$), and the beacon position ${}^{\mathcal{I}}\bm{r}_{\rm{pl}}$ can be defined as
@@
${\delta}\ {}^{\mathbb{C}}\bm{r}_{\rm{pl}}=\underbrace{\begin{bmatrix}\dfrac{\partial\ {}^{\mathbb{C}}\bm{r}_{\rm{pl}}}{\partial q_{0}}&\Big{[}\dfrac{\partial\ {}^{\mathbb{C}}\bm{r}_{\rm{pl}}}{\partial\bm{q}_{v}}\Big{]}&\Big{[}\dfrac{\partial\ {}^{\mathbb{C}}\bm{r}_{\rm{pl}}}{\partial^{\mathcal{I}}\bm{r}}\Big{]}&\Big{[}\dfrac{\partial\ {}^{\mathbb{C}}\bm{r}_{\rm{pl}}}{\partial^{\mathcal{I}}\bm{r}_{\rm{pl}}}\Big{]}\\ \end{bmatrix}}_{\bm{G}}\begin{pmatrix}\delta q_{0}\\ \delta\bm{q}_{v}\\ \delta^{\mathcal{I}}\bm{r}\\ \delta^{\mathcal{I}}\bm{r}_{\rm{pl}}\\ \end{pmatrix}$ (12)
Here the quaternion $\bm{q}_{\mathcal{C}/\mathcal{I}}=\left(q_{0},\,\bm{q}_{v}\right)^{\top}$ represents the rotation from inertial reference frame $\mathcal{I}$ to the camera reference frame $\mathcal{C}$. The matrix $\bm{G}$ has dimension $2\times 10$ and it can be expressed as
@@
$\bm{G}=\underbrace{\begin{bmatrix}\dfrac{1}{{}_{h}^{\mathbb{C}}{r}_{\rm{pl}_{3}}}&0&-\dfrac{{}_{h}^{\mathbb{C}}{r}_{\rm{pl},(1)}}{{}_{h}^{\mathbb{C}}{r}_{\rm{pl},(3)}^{2}}\\ 0&\dfrac{1}{{}_{h}^{\mathbb{C}}{r}_{\rm{pl},(3)}}&-\dfrac{{}_{h}^{\mathbb{C}}{r}_{\rm{pl},(2)}}{{}_{h}^{\mathbb{C}}{r}_{\rm{pl},(3)}^{2}}\end{bmatrix}}_{\bm{D}}\bm{K}_{\rm{cam}}\begin{bmatrix}\dfrac{\partial{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}^{\mathcal{C}}r_{\textrm{pl}}}}{\partial q_{0}}&\dfrac{\partial{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}^{\mathcal{C}}r_{\textrm{pl}}}}{\partial\bm{q}_{v}}&\dfrac{\partial{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}^{\mathcal{C}}r_{\textrm{pl}}}}{\partial^{\mathcal{I}}\bm{r}}&\dfrac{\partial{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}^{\mathcal{C}}r_{\textrm{pl}}}}{\partial^{\mathcal{I}}\bm{r}_{\rm{pl}}}\end{bmatrix}$ (13)
@@
with $\bm{D}$ the matrix containing the derivatives of ${}^{\mathbb{C}}{\bm{r}}_{\rm{pl}}$ with respect to ${}_{h}^{\mathbb{C}}{\bm{r}}_{\rm{pl}}$, and ${}_{h}^{\mathbb{C}}{r}_{\rm{pl},(i)}$ the $i$-esimal component of the object location in homogenous coordinates. Note that the partial derivatives of ${}^{\mathcal{C}}r_{\textrm{pl}}=\bm{A}_{\textrm{corr}}\ (^{\mathcal{I}}\bm{r}_{\rm{pl}}-^{\mathcal{I}}\bm{r})$ with respect to the spacecraft pose and beacon position are
$\frac{\partial{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}^{\mathcal{C}}r_{\textrm{pl}}}}{\partial q_{0}}=2q_{0}{}^{\mathcal{I}}{\bm{\rho}}-2[\bm{q}_{v}]^{\wedge}\ {}^{\mathcal{I}}{\bm{\rho}}$ (14)
$\frac{\partial{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}^{\mathcal{C}}r_{\textrm{pl}}}}{\partial\bm{q}_{v}}=-2\ {}^{\mathcal{I}}{\bm{\rho}}\bm{q}_{v}^{\top}+2\bm{q}_{v}^{\top}\ {}^{\mathcal{I}}{\bm{\rho}}\bm{I}_{3x3}+2\bm{q}_{v}\ {}^{\mathcal{I}}{\bm{\rho}^{\top}}+2q_{0}[{}^{\mathcal{I}}{\bm{\rho}}]^{\wedge}$ (15)
$\frac{\partial{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}^{\mathcal{C}}r_{\textrm{pl}}}}{\partial^{\mathcal{I}}\bm{r}}=-\bm{A}_{\textrm{corr}}$ (16)
$\frac{\partial{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}^{\mathcal{C}}r_{\textrm{pl}}}}{\partial^{\mathcal{I}}\bm{r}_{{\rm{pl}}}}=\bm{A}_{\textrm{corr}}$ (17)
@@
$\delta q_{0}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}1}$ (18)
$\delta\bm{q}_{v}=\frac{1}{2}\delta(\theta\bm{e})$ (19)
It is worth noting that the variance of the scalar part is $\sigma_{q_{0}}^{2}=0$. Therefore, $\bm{S}$ can be described:
$\bm{S}=\rm{diag}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{q_{0}}^{2}},\sigma_{\bm{q}_{v}}^{2}{\bm{I}_{3x3}},\sigma_{\bm{r}}^{2}{\bm{I}_{3x3}},\sigma_{\bm{r}_{\rm{pl}}}^{2}{\bm{I}_{3x3}})$ (20)
@@
$a=\sqrt{11.8292\ \lambda_{\text{max}}}$ $b=\sqrt{11.8292\ \lambda_{\text{min}}}$ $\psi={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\arctan{\Big{(}\dfrac{v_{\rm{max}_{2}}}{v_{\rm{max}_{1}}}\Big{)}}}$ ) (21)
@@
4 Extended Kalman Filter Based on Planets Observations
In this section, the VBN filter is described. Firstly, the dynamic and measurement models adopted in the VBN filter are detailed. Successively, the chosen filtering scheme is shown. Note that the vectors specified in this section are always defined in the inertial reference frame $\mathcal{I}$. Thus, the superscript is indicated only for exceptions.
@@ 4.1 Dynamics Model
The dynamics model adopted in this work is an extension of the one described in Andreis et al. [9]. The process state $\bm{x}$ is defined as
@@
where ${}^{\mathcal{}}\bm{r}$ and ${}^{\mathcal{}}\bm{v}$ are the inertial probe position and velocity, respectively, and $\bm{\eta}=[\bm{\eta}_{\textrm{R}},\bm{\eta}_{\textrm{SRP}}]^{\top}$ is a vector of Gauss–Markov (GM) processes accounting for unmodeled terms: a 3-dimensional residual accelerations $\bm{\eta}_{\textrm{R}}$ and the stochastic component of the Solar Radiation Pressure (SRP) $\bm{\eta}_{\textrm{SRP}}$ [40]. The process is modeled using the following differential equation
@@
$\dot{\bm{x}}(t)=\underbrace{\begin{bmatrix}&&{}^{\mathcal{}}\bm{v}&&\\ &&{}^{\mathcal{}}\bm{a}_{\rm{Sun}}+{}^{\mathcal{}}\bm{a}_{\rm{SRP}}+\sum_{i}{}^{\mathcal{}}\bm{a}_{\rm{pl}_{i}}+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{\eta}_{\textrm{R}}+\bm{\eta}_{\textrm{SRP}}}&&\\ &&-\xi_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textrm{R}}}\bm{\eta}_{\textrm{R}}&&\\ &&-\xi_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textrm{SRP}}}\bm{\eta}_{\textrm{SRP}}&&\\ \end{bmatrix}}_{\bm{f}}+\underbrace{\begin{bmatrix}&&\bm{0}_{3x1}&&\\ &&{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{0}_{3x1}}&&\\ &&\bm{w}_{\textrm{R}}&&\\ &&\bm{w}_{\textrm{SRP}}&&\\ \end{bmatrix}}_{\bm{w}}$ (24)
@@
As detailed in Jean et al. [41], the terms that describe the SRP are: $C_{\textrm{R}}$ the coefficient of reflection, $P_{\textrm{0}}$ the solar power, $R_{\textrm{0}}$ the Sun radius, $A_{\textrm{s}}$ the cross-section area of the probe, and $m_{\textrm{s}}$ its mass. The third-body perturbations of the Earth-Moon barycenter, Mars, and Jupiter are also included. Recall that $\mu_{\rm{pl}_{i}}$ is the gravitational parameter of the $i$-esimal planet considered. In the Langevin equations that govern the Gauss-Markov processes the coefficients $\xi_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textrm{R}}}$ and $\xi_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textrm{SRP}}}$ define the reciprocal of the correlation times, while $\bm{w}_{\textrm{R}}$ and $\bm{w}_{\textrm{SRP}}$ are the process noises of the Gauss-Markov parameters with $\sigma_{\textrm{R}}$ and $\sigma_{\textrm{SRP}}$ standard deviations. The covariance of these two process noises can be defined as
$\textrm{E}[\bm{w}_{\textrm{R}}\bm{w}_{\textrm{R}}^{\top}]=\sigma_{\textrm{R}}^{2}\bm{I}_{3x3}=\bm{Q}_{\rm{R}}\qquad\textrm{E}[\bm{w}_{\textrm{SRP}}\bm{w}_{\textrm{SRP}}^{\top}]=\sigma_{\textrm{SRP}}^{2}\bm{I}_{3x3}=\bm{Q}_{\rm{SRP}}$ (28)
These covariances are not apriori known. Yet, reference values for the acceleration uncertainties $Q_{\rm{a_{R}}}$ and $Q_{\rm{a_{SRP}}}$ are reported in the literature. Moreover, these covariances $Q_{\rm{a_{R}}}$ and $Q_{\rm{a_{SRP}}}$ can be treated as steady-state covariance values of the GM parameters [40]. That is,
$\textrm{E}[\bm{\eta}_{\rm{R}}\bm{\eta}_{\rm{R}}^{\top}]=\frac{\bm{Q}_{\rm{R}}}{2\xi_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textrm{R}}}}=\bm{Q}_{\rm{a_{R}}}\qquad\rm{E}[\bm{\eta}_{\textrm{SRP}}\bm{\eta}_{\rm{SRP}}^{\top}]=\frac{\bm{Q}_{\rm{SRP}}}{2\xi_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textrm{SRP}}}}=\bm{Q}_{\rm{a_{SRP}}}$ (30)
From which $\bm{Q}_{\rm{{SRP}}}$ and $\bm{Q}_{\rm{{R}}}$ can be computed. The complete process noise covariance matrix is $\bm{Q}$:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{Q}=\rm{diag}(\bm{0}_{3x3},\bm{0}_{3x3},\bm{Q}_{\rm{R}},\bm{Q}_{\rm{SRP}})}$ (31)

### v3

@@ 3.2 Light-Aberration Correction
After the first attitude determination, the centroids of the stars are corrected for the light-aberration effect, and the probe attitude is recomputed by taking into account the corrected stars LoS directions. The procedure adopted is described in [19]. At first, the observed stars LoS directions as seen by the spacecraft in the inertial reference frame $\mathcal{N}$ are found:
${}^{\mathcal{N}}\bm{\rho}_{s_{\rm{obs}}}=(\bm{K}_{\rm{cam}}\bm{A})^{-1}{}_{h}^{\mathbb{C}}\bm{r}_{s_{\rm{obs}}}$ (3)
where ${}_{h}^{\mathbb{C}}\bm{r}_{s_{\rm{obs}}}$ are the observed stars position projections in $\mathbb{C}$ in homogeneous coordinates (see [35] for homogeneous coordinates), and $\bm{K}_{\rm{cam}}$ is the camera calibration matrix. Then, the angle $\theta_{\textrm{obs}}$ between the observed unitary stars LoS directions ${}^{\mathcal{N}}\hat{\bm{\rho}}_{s_{\rm{obs}}}$ and the estimated unitary velocity vector of the probe $\hat{\bm{v}}_{\rm{p}}$ is defined as
$\tan{\theta_{\textrm{obs}}}=\frac{||{}^{\mathcal{N}}\hat{\bm{\rho}}_{s_{\rm{obs}}}\times\hat{\bm{v}}_{\rm{p}}||}{{}^{\mathcal{N}}\hat{\bm{\rho}}_{s_{\rm{obs}}}^{\top}\hat{\bm{v}}_{\rm{p}}}$ (4)
And the aberration angle $\varepsilon$ is evaluated:
$\tan{\varepsilon}=\dfrac{(v_{\rm{p}}/c)\sin\theta_{\textrm{obs}}}{1-(v_{\rm{p}}/c)\cos\theta_{\textrm{obs}}}$ (5)
where $c$ is the speed of the light. Thus, the corrected unitary stars LoS directions ${}^{\mathcal{N}}\hat{\bm{\rho}}_{s_{\rm{corr}}}$ can be retrieved as
${}^{\mathcal{N}}\hat{\bm{\rho}}_{s_{\rm{corr}}}=\dfrac{{}^{\mathcal{N}}\hat{\bm{\rho}}_{s_{\textrm{obs}}}\sin\theta_{\textrm{corr}}-\hat{\bm{v}}_{\rm{p}}\sin\varepsilon}{\sin{\theta_{\textrm{obs}}}}$ (6)
with $\theta_{\textrm{corr}}=\theta_{\textrm{obs}}+\varepsilon$. At this point, the attitude matrix of the probe is redetermined by solving the Wahba problem [33] in which ${}^{\mathcal{N}}\hat{\bm{\rho}}_{s_{\rm{corr}}}$ are considered. This corrected attitude matrix is labeled $\bm{A}_{\rm{corr}}$.
@@ 3.3 Beacon Identification
At this step, the planet must be identified in the image and its projection ${}^{\mathbb{C}}\bm{r}_{\rm{pl}}$ must be extracted. In this step, the identification is performed through the evaluation of the statistical moments associated with the planet position projection, which defines the Gaussian probability of finding the planet in that portion of the image. At first, the expected position projection of the observed planet is evaluated as:
${}_{h}^{\mathbb{C}}\bm{r}_{\rm{pl}_{0}}=\bm{K}_{\rm{cam}}\bm{A}_{\rm{corr}}(^{\mathcal{N}}\bm{r}_{\rm{pl}}-^{\mathcal{N}}\bm{r}_{\rm{p}})$ (7)
@@
The covariance matrix of the beacon position projection $\bm{P}$ due to the spacecraft pose and beacon position uncertainty is computed as
@@
where $\bm{G}$ is the Jacobian matrix of the mapping between the beacon position projection ${}^{\mathbb{C}}\bm{r}_{\rm{pl}}$ and the spacecraft pose and the beacon position, and $\bm{S}$ is the uncertainty covariance matrix of the probe pose and beacon position. To evaluate $\bm{G}$, the variation of $\ {}^{\mathbb{C}}\bm{r}_{\rm{pl}}$ with respect to the variation of the spacecraft pose and the beacon position is computed. In particular, the quaternions $\bm{q}=(q_{0},\bm{q}_{v})^{\top}$ are chosen to represent the probe attitude matrix. Eq. (9) gives the quaternion representation of the attitude matrix $\bm{A}_{\textrm{corr}}$ [33]
@@
where $\left[{(\cdot)}\right]^{\wedge}$ is the skew-symmetric matrix associated with the cross-product operation. Thus, the variation of $\ {}^{\mathbb{C}}\bm{r}_{\rm{pl}}$ with respect to the variation of the spacecraft pose, i.e., $\bm{A}_{\textrm{corr}}(\bm{q}_{\mathcal{C}/\mathcal{N}})$ and ${}^{\mathcal{N}}\bm{r}$, and the beacon position ${}^{\mathcal{N}}\bm{r}_{\rm{pl}}$ can be defined as
@@
${\delta}\ {}^{\mathbb{C}}\bm{r}_{\rm{pl}}=\underbrace{\begin{bmatrix}\dfrac{\partial\ {}^{\mathbb{C}}\bm{r}_{\rm{pl}}}{\partial q_{0}}&\Big{[}\dfrac{\partial\ {}^{\mathbb{C}}\bm{r}_{\rm{pl}}}{\partial\bm{q}_{v}}\Big{]}&\Big{[}\dfrac{\partial\ {}^{\mathbb{C}}\bm{r}_{\rm{pl}}}{\partial^{\mathcal{N}}\bm{r}}\Big{]}&\Big{[}\dfrac{\partial\ {}^{\mathbb{C}}\bm{r}_{\rm{pl}}}{\partial^{\mathcal{N}}\bm{r}_{\rm{pl}}}\Big{]}\\ \end{bmatrix}}_{\bm{G}}\begin{pmatrix}\delta q_{0}\\ \delta\bm{q}_{v}\\ \delta^{\mathcal{N}}\bm{r}\\ \delta^{\mathcal{N}}\bm{r}_{\rm{pl}}\\ \end{pmatrix}$ (10)
The matrix $\bm{G}$ has dimension $2\times 10$ and it is defined as
@@
$\bm{G}={\begin{bmatrix}\dfrac{1}{{}_{h}^{\mathbb{C}}{r}_{\rm{pl}_{3}}}&0&-\dfrac{{}_{h}^{\mathbb{C}}{r}_{\rm{pl}_{1}}}{{}_{h}^{\mathbb{C}}{r}_{\rm{pl}_{3}}^{2}}\\ 0&\dfrac{1}{{}_{h}^{\mathbb{C}}{r}_{\rm{pl}_{3}}}&-\dfrac{{}_{h}^{\mathbb{C}}{r}_{\rm{pl}_{2}}}{{}_{h}^{\mathbb{C}}{r}_{\rm{pl}_{3}}^{2}}\end{bmatrix}}\bm{K}_{\rm{cam}}\begin{bmatrix}\dfrac{\partial(\bm{A}_{\textrm{corr}}\ {}^{\mathcal{N}}{\bm{\rho}})}{\partial q_{0}}&\dfrac{\partial(\bm{A}_{\textrm{corr}}\ {}^{\mathcal{N}}{\bm{\rho}})}{\partial\bm{q}_{v}}&\dfrac{\partial(\bm{A}_{\textrm{corr}}\ {}^{\mathcal{N}}{\bm{\rho}})}{\partial^{\mathcal{N}}\bm{r}}&\dfrac{\partial(\bm{A}_{\textrm{corr}}\ {}^{\mathcal{N}}{\bm{\rho}})}{\partial^{\mathcal{N}}\bm{r}_{\rm{pl}}}\end{bmatrix}$ (11)
@@
where the partial derivatives of $(\bm{A}_{\textrm{corr}}\ {}^{\mathcal{N}}{\bm{\rho}})$ with respect to the spacecraft pose and beacon position are
$\frac{\partial(\bm{A}_{\textrm{corr}}\ {}^{\mathcal{N}}{\bm{\rho}})}{\partial q_{0}}=2q_{0}{}^{\mathcal{N}}{\bm{\rho}}-2[\bm{q}_{v}]^{\wedge}\ {}^{\mathcal{N}}{\bm{\rho}}$ (12)
$\frac{\partial(\bm{A}_{\textrm{corr}}\ {}^{\mathcal{N}}{\bm{\rho}})}{\partial\bm{q}_{v}}=-2\ {}^{\mathcal{N}}{\bm{\rho}}\bm{q}_{v}^{\top}+2\bm{q}_{v}^{\top}\ {}^{\mathcal{N}}{\bm{\rho}}\bm{I}_{3x3}+2\bm{q}_{v}\ {}^{\mathcal{N}}{\bm{\rho}^{\top}}+2q_{0}[{}^{\mathcal{N}}{\bm{\rho}}]^{\wedge}$ (13)
$\frac{\partial(\bm{A}_{\textrm{corr}}\ {}^{\mathcal{N}}{\bm{\rho}})}{\partial^{\mathcal{N}}\bm{r}}=-\bm{A}_{\textrm{corr}}$ (14)
$\frac{\partial(\bm{A}_{\textrm{corr}}\ {}^{\mathcal{N}}{\bm{\rho}})}{\partial^{\mathcal{N}}\bm{r}_{{\rm{pl}}}}=\bm{A}_{\textrm{corr}}$ (15)
@@
$\delta q_{0}=0$ $\delta\bm{q}_{v}=\frac{1}{2}\delta(\theta\bm{e})$ (16)
Since $\sigma_{\bm{q}_{0}}^{2}=0$ for the small-angles assumption, $\bm{S}$ can be described:
$\bm{S}=\rm{diag}(\sigma_{\bm{q}_{v}}^{2}{\bm{I}_{3x3}},\sigma_{\bm{r}}^{2}{\bm{I}_{3x3}},\sigma_{\bm{r}_{\rm{pl}}}^{2}{\bm{I}_{3x3}})$ (17)
@@
$a=\sqrt{11.8292\ \lambda_{\text{max}}}$ $b=\sqrt{11.8292\ \lambda_{\text{min}}}$ $\psi=\arctan{\Big{(}\frac{\bm{v}_{\rm{max}_{2}}}{\bm{v}_{\rm{max}_{1}}}\Big{)}}$ ) (18)
@@
4 Non-dimensional Extended Kalman Filter Based on Planets Observations
In this section, the VBN filter is described. Firstly, the dynamic and measurement models adopted in the VBN filter are detailed. Successively, the chosen filtering scheme is shown. Note that the vectors specified in this section are always defined in the inertial reference frame $\mathcal{N}$. Thus, the superscript is indicated only for exceptions.
@@ 4.1 Dynamics Model
The process state $\bm{x}$ is defined as
@@
where ${}^{\mathcal{}}\bm{r}$ and ${}^{\mathcal{}}\bm{v}$ are the inertial probe position and velocity, respectively, and $\bm{\eta}$ is a vector of Gauss–Markow (GM) processes accounting for unmodeled terms: a 3-dimensional residual accelerations $\bm{\eta}_{\textrm{R}}$ and the stochastic component of the Solar Radiation Pressure (SRP) $\bm{\eta}_{\textrm{SRP}}$; that is, $\bm{\eta}=[\bm{\eta}_{\textrm{R}},\bm{\eta}_{\textrm{SRP}}]^{\top}$ [36]. The process is modeled using the following equation of motion
@@
$\dot{\bm{x}}(t)=\underbrace{\begin{bmatrix}&&{}^{\mathcal{}}\bm{v}&&\\ &&{}^{\mathcal{}}\bm{a}_{\rm{Sun}}+{}^{\mathcal{}}\bm{a}_{\rm{SRP}}+\sum_{i}{}^{\mathcal{}}\bm{a}_{\rm{pl}_{i}}&&\\ &&-\xi\bm{\eta}_{\textrm{R}}&&\\ &&-\xi\bm{\eta}_{\textrm{SRP}}&&\\ \end{bmatrix}}_{\bm{f}}+\underbrace{\begin{bmatrix}&&\bm{0}_{3x1}&&\\ &&\bm{\eta}_{\textrm{R}}+\bm{\eta}_{\textrm{SRP}}&&\\ &&\bm{w}_{\textrm{R}}&&\\ &&\bm{w}_{\textrm{SRP}}&&\\ \end{bmatrix}}_{\bm{w}}$ (21)
@@
The terms that describe the SRP are [37]: $C_{\textrm{R}}$ the coefficient of reflection, $P_{\textrm{0}}$ the solar power, $R_{\textrm{0}}$ the Sun radius, $A_{\textrm{s}}$ the cross-section area of the probe, and $m_{\textrm{s}}$ its mass. The third-body perturbations of the Earth-Moon barycenter, Mars, and Jupiter are included; $\mu_{\rm{pl}_{i}}$ is the gravitational parameter of the $i$-esimal planet considered. In the Langevin equations that govern the GM processes the coefficient $\xi$ defines the reciprocal of the correlation time, while $\bm{w}_{\textrm{R}}$ and $\bm{w}_{\textrm{SRP}}$ are the process noises of the GM parameters with $\sigma_{\textrm{R}}$ and $\sigma_{\textrm{SRP}}$ standard deviations, respectively [36]. The process noise covariance matrix is $\bm{Q}$:
$\bm{Q}=\textrm{diag}(\bm{0}_{3x3},\bm{Q}_{a},\bm{Q}_{\rm R},\bm{Q}_{\rm SRP})$ (25)
with $\bm{Q}_{\textrm{R}}=\sigma_{\textrm{R}}^{2}\bm{I}_{3x3}$, $\bm{Q}_{\textrm{SRP}}=\sigma_{\textrm{SRP}}^{2}\bm{I}_{3x3}$, and $\bm{Q}_{\rm a}=(\bm{Q}_{\rm R}+\bm{Q}_{\rm SRP})/(2\xi)$.